OpenStudy (anonymous):

Simplify using properties : 2(a + 4) + -2a + -9

OpenStudy (anonymous):

\[2(a + 4) - 2a - 9\] Which properties?

OpenStudy (anonymous):

2a+8-2a-9 =-1

OpenStudy (anonymous):

Using distributive property, 2(a + 4) - 2a - 9 = 2a + 8 - 2a - 9 Using commutative property, 2a - 2a + 8 - 9 = 0 + -1 = -1

OpenStudy (anonymous):

@JessCALI welcome to OpenStudy(OS)!!!!!!!!